A silver necklace originaly priced at $45.50 is on sale for 1/5 off. what is the sale price?
i can't understand how to to calculate this. please show me in detail.
thanku!

Found 2 solutions by KBORNSGIRL, Linz:
Answer by KBORNSGIRL(1)

Your silver necklace is worth $45.50 and the sale price is for 1/5 off. First you should find out how much is 1/5.
If you divide 5 into 1 you will get .2 this is your percentage off.
Then you take your silver necklace price of $45.50 and multiply it by your percentage of .2 which will give you $9.10.
$45.50 - $9.10 = $36.40 is what you would pay for this silver necklace.

Answer by Linz(9)

1/5 is = to 20%
so, you multiply 45.50 by 20%
45.50*20%=9.10
Then you subract 9.10 from 45.50
45.50-9.10=36.40
So the sale price is $36.40